### 08 Dec 2023 v1

# Filtering Pixel Latent Variables for Unmixing Volumetric Images

Abstract
Measurements of different overlapping components require robust unmixing algorithms to convert the raw multi-dimensional measurements to useful unmixed images. Such algorithms perform reliable separation of the components when the raw signal is fully resolved and contains enough information to fit curves on the raw distributions. In experimental physics, measurements are often noisy, undersampled, or unresolved spatially or spectrally. We propose a novel method where bandpass filters are applied to the latent space of a multi-dimensional convolutional neural network to separate the overlapping signal components and extract each of their relative contributions. Simultaneously processing all dimensions with multi-dimensional convolution kernels empowers the network to combine the information from adjacent pixels and time- or spectral-bins, facilitating component separation in instances where individual pixels lack well-resolved information. We demonstrate the applicability of the method to real experimental physics problems using fluorescence lifetime microscopy and mode decomposition in optical fibers as test cases. The successful application of our approach to these two distinct experimental cases, characterized by different measured distributions, highlights the versatility of our approach in addressing a wide array of imaging tasks.
1 Introduction
Unmixing refers to the technique of breaking down a measurement into its unique constituents and their respective proportions. This concept is commonly linked to spectral analysis, where the overlapping spectral signatures of different emitters hinder the accurate identification of each emitter’s abundance, especially when the spectral signature is undetermined or the data is insufficiently sampled. Beyond spectral measurements, unmixing applies to any measurements that combine spatial dimensions (x,y,z) to additional continuous and often multi-channel measures. In fluorescence lifetime microscopy, this additional dimension is the time it took for the photon to be emitted[4]. In spectrally and spatially (S) resolved measurements for characterizing optical fibers, this dimension is the amplitude of the propagated light for different input wavelengths [21].
In this work, we propose a generalized unmixing method based on convolutional neural networks that applies to various types of multi-dimensional measurements (Fig. 1). We specifically focus on noisy data or undersampled distributions, for which common methods fail to accurately unmix the different contributions. The proposed method relies on transforming the overlapping input distributions to a latent space where they are untangled and can be separated using simple bandpass filters. We demonstrate the method on synthetic datasets based on real data and created for real experimental physics use cases, showcasing the general applicability of the approach.
2 Related work
In this work, we demonstrate an unmixing method on two experimental physics use cases: fluorescence lifetime microscopy (FLIM) and mode decomposition in optical fibers. In both cases, the input data is composed of overlapping distributions that need to be decomposed into single channel outputs to retrieve useful information.
2.1 Multidimensional CNNs and the latent space
Three-dimensional Convolutional Neural Networks (3D CNNs) offer advantages over their two-dimensional counterparts in tasks involving volumetric imaging such as medical imaging [28] or spatiotemporal data such as videos [14]. Their ability to capture both spatial and temporal features simultaneously leads to improved representations of complex patterns in three-dimensional space [32]. Unmixing tasks face the challenge of combining 3D input with 2D multi-channels output, making the task difficult to adapt to existing 3D models; for example, with a 3D U-Net, where the 3D features learned in the encoder cannot be directly linked to the 2D features required in the decoder. CNNs that take as input a 3D volume and predict a 2D image is not a common architecture as it is a distinct task from volumetric segmentation, and no architecture has been developed specifically for this task. Tasks on 3D volumes are generally either in the volumetric space from input to output (3D segmentation [28]) or a classification layer is added to the learned representation for a 1D output (classification [27]). 2.5D networks have been proposed for the segmentation of volumes into 2D slice-wise segmentation for cases where 3D CNNs are not applicable because of slice discontinuities [23], or in cases where the data is too limited to train 3D networks [15]. However, 2.5D networks were shown to be less accurate than 3D CNNs when 3D CNNs are applicable [2].
In our method, the features are represented in the 3D space from the entry to the exit of the network, so that existing 3D architectures can be exploited for unmixing tasks. This leads to a general model that is not constrained in the number of target output channels, because the learned features are translated into channels directly from the learned 3D latent space. This latent space is an abstract and simplified representation of the data; in convolutional neural networks, the latent space simplifies the raw input pixel data into an informative representation that is used by the subsequent layers to perform the task[17].
2.2 Fluorescence lifetime microscopy
Fluorescence lifetime microscopy (FLIM) uses the arrival time of fluorescence photons to assign them to specific fluorophores or to measure environmental changes in biological samples [4]. The fluorescence decay follows an exponential probability distribution characterized by the parameter corresponding to the fluorescence lifetime [5]. The fluorescence lifetime represents a distinctive photophysical characteristic unique to each fluorescent molecule, which is used in FLIM to discriminate between several fluorophores. However, even by carefully selecting molecules based on their fluorescence lifetime, exponential distributions always overlap and signal unmixing can be challenging, especially in the low photon-count regime (100 photons/pixel) or when multiple fluorophores are present at a given position. For live-cell imaging, where photobleaching and sample movement should be minimized, FLIM measurements are generally associated with low photon-counts [31]. Additionally, measurements of interactions between multiple components increase the complexity of assigning each photon to a given emitting molecule.
2.2.1 FLIM analysis
Different unmixing approaches have been proposed for FLIM such as Maximum Likelihood Estimation (MLE) [3], phasor analysis [8], and deep learning based strategies [30]. MLE is an effective method for fitting parameters to describe a curve given a set of samples from the distribution and is a common approach for FLIM data analysis [3]. An alternative approach relying on phasor analysis has been proposed for unmixing fluorescence lifetime distributions [8]. The performance of both approaches is strongly limited by low photon counts (300) and when more than three components have to be unmixed [7, 6, 25, 8].
Recently, deep learning based approaches have been proposed for FLIM data analysis. The Fluorescence lifetime imaging network (FLI-Net) [30] is a 3D-CNN that takes as input a 3D FLIM volume (x,y,t) where each pixel is characterized by photons sampled in a biexponential distribution with aim to assign a mean lifetime to each pixel rather than unmixing the input image. The Fluorescence Lifetime AdderNet is a 1D network trained for lifetime unmixing [33], which can handle spatially distinct structures and does not use the spatial features to improve photon assignment in overlapping structures.
2.2.2 Unmixing strategies for fluorescence microscopy
A common method for unmixing multi-channel images from a single acquisition is to combine the lifetime and spectral information (spectral FLIM, S-FLIM). In this case, acquired volumes are 4D (x,y,,t). The multidimensional information facilitates the assignment of photons to a given fluorescent molecule, since emitters characterized by similar lifetimes can have different spectral signatures or vice versa. Phasor analysis has been explored as a method for unmixing S-FLIM data [19, 10]. Deep learning-based unmixing of S-FLIM data has also been demonstrated with UNMIX-ME [29], which uses convolutional neural networks to map noisy hyperspectral FLIM data to spatially independent unmixed fluorescence coefficient values. The method allows the separation of up to four components in simulated samples.
To our knowledge, there is a gap in the literature for FLIM unmixing method that applies to images in the low photon count regime (100 photon/pixel) and more than two fluorescent emitters per pixel. Such an approach would be transformative for the democratization of FLIM for multi-channel live-cell fluorescence microscopy and nanoscopy.
2.3 Mode decomposition in multi-mode fibers
Multi-mode fibers are optical fibers that support the propagation of multiple simultaneous light modes [11]. Spatially and spectrally resolved (S) imaging[21] is a measurement commonly used for characterizing how light travels through multi-mode optical fibers. By measuring the spatial distribution of the intensity at the output of the fiber for a range of laser wavelengths, the modes of propagation of the light through the fiber can be characterized. The 2D image at the fiber output is not merely a linear combination of each distinct mode because the coherent modes interfere constructively or destructively with each other in function of their relative phases. In a two modes fiber, the intensity of a single pixel versus the wavelength of the laser propagating through the fiber, within a small spectral range, oscillates periodically with a sinusoidal distribution (Fig. 7). When more than one mode propagates through the fiber, the intensity oscillates as the sum of each sinus function associated to each combination of two modes.
In a properly sampled measurement where the frequency of acquisition is above the Nyquist frequency, these oscillations allow to separate each mode with a Fourier transform based method [20] or by principal component analysis (PCA) [24, 26]. Other methods exist to achieve real-time mode decomposition from a single image with various algorithms [13, 1, 18], but they all require the a priori knowledge of the spatial distribution of each mode.
To our knowledge, no method for mode decomposition in multi-mode fibers solve both limitations of current approaches: separate modes with unknown spatial distributions, and successfully retrieve the contribution of each mode with a sub-Nyquist sampling rate.
3 Methodology
3.1 Datasets
Experiments are conducted with three datasets: 1) a modified version of the MNIST dataset [16] with overlaid digits and simulated lifetime (3D-MNIST), 2) a FLIM dataset generated from STimulated Emission Depletion (STED) microscopy images of neuronal proteins (FLIM-STED), and 3) a simulation of spectrally and spatially resolved measurements of light propagating in multi-mode fibers (S-MMF).
3D-MNIST The 3D-MNIST dataset was used to develop the method and evaluate the influence of different parameters. Images from the classes 4, 8 and 9 are overlaid to create mixed images. Random Poisson noise is generated over the whole image as a fourth target channel. This multi-dimensional version of MNIST is generated according to Algo. 1.
The generated MNIST dataset is parametrized by a signal-to-noise factor , decay constants for each channel (, , , ), a time dimension length , and a number of images . Each data sample is created by randomly sampling up to three digits from classes (4, 8, 9) from the MNIST training set. We chose the digits 4, 8 and 9 because this triplet of digits has the highest fraction of overlap (Supp. fig. 2). There is a probability 0.5 of picking one digit for each of the (4,8,9) classes (lines 9 to 13), such that a training sample can contain 0, 1, 2 or 3 digits. The images are independently randomly rotated, flipped and translated (line 17). To reduce the pixel counts from [0,255] to a noisier range more representative of real experimental conditions, the image is first normalized to the range [0,1] by dividing each image by its maximum value (line 19). A random count is assigned to each pixel by sampling an intensity value in a Poisson distribution parameterized by the pixel value, multiplied by the SNR factor (lines 21 and 22). For each count, an equivalent number of time values are sampled following an exponential distribution parameterized by the decay constant of the channel (line 24). The time values of each individual pixel are converted to a histogram of bins of fixed width (line 26). A background channel is also created by randomly sampling values in a Poisson distribution with parameter and creating a time histogram following the same method as for the digits channels (lines 29 to 33). The four volumes (4, 8, 9, background) are added along the channel dimension to create a mixed volume of dimensions () (line 35). The process is repeated times to generate the 3D-MNIST dataset.
The 3D-MNIST dataset counts 8,000 combined images generated with and , from which 6,400 randomly sampled images were used for training and 1,600 for testing.
STED-FLIM The STED-FLIM dataset consists of real microscopy images of four different proteins tagged with fluorescent markers in fixed cortical neurons (Suppl. sec. A). The signal from each fluorescent marker was spectrally separated to obtain ground truth images with four independent channels. The STED-FLIM dataset was generated by adapting the 3D-MNIST data generation algorithm (algo. 1). The following modifications were applied to the algorithm: 1) dataset is a set of 4-color FLIM images, 2) channels are not randomly discarded and independently sampled (lines 10-15), 3) the true measured photon counts were preserved instead of normalized and randomized (lines 16-22), 4) no background channel was added (lines 29-33). The lifetime values (Suppl. table 1) were chosen to fit within the range of commercially available fluorescent dyes for STED microscopy [12]. We chose values within this range so that the separation between lifetimes is variable (0.7 ns A-B, 0.6 ns B-C, 0.2 ns C-D). Since the lifetime distributions are synthetic, the actual lifetimes do not need to match a specific dye, so we attributed generic names (A, B, C and D) to the fluorescent markers for clarity. Channels A and C were acquired with confocal microscopy while channels B and D were acquired with STED microscopy. This allows to test if the method can adapt to varying photon counts and spatial resolutions. The STED-FLIM dataset counts 400 training and validation images which were preprocessed into 8,261 crops, and 20 test images.
| Protein | Marker | Lifetime [ns] |
| MAP2 | A | 2.6 |
| PSD95 | B | 3.3 |
| VGLUT1 | C | 3.9 |
| FUS | D | 4.1 |
S-MMF The S-MMF dataset consists in images of the intensity measured at the output of optical fibers for different input wavelengths [22]. The simulated measure has three dimensions; two spatial (2D images of the light distribution at the fiber output) and one spectral (each image represents the measure obtained from a different light wavelength). The dataset was computed with the equations described in Quiquempois et al. (2021) [24]. A vector finite difference mode-solver [9] was used to find mode properties around 1550 nm for various index profiles and fiber geometries. All fibers in the dataset support up to 6 modes. The mode spatial distribution and effective index dependency on wavelength (dispersion) were taken into account. For all data samples, we considered a spectral range of 6 nm for a fiber length of 5 m. The dataset includes measurements with various combinations of up to 6 propagating modes, four different fiber geometries (step-index with varying numerical apertures and core size, GRIN, W-type, H-type and a realistic step-index geometry with a central dip), and different mode-specific intensities (). We used 1500 fiber measurements to train the network and 175 for testing.
3.2 Model
Network architecture 3D convolution kernels allow the processing of all dimensions of a 3D input volume simultaneously. We chose to use a 3D CNN for unmixing undersampled data so that the 3D kernels can combine essential information from neighboring pixels and neighboring time- or spectral-bins. The network used here is a standard 3D U-Net (Supp. sec. B). As is standard for U-Nets, the dimensionality of the input remains unchanged through the layers; the input 3D volume results in a 3D output of the same size. We added a layer with one convolution filter at the end of the network that maps the output of the U-Net to a single-channel volumetric output of dimension , with and coming from the images spatial dimensions and being the pixel latent space size.
Latent space filters For unmixing applications, the input is a 3D volume and the target is a channel 2D image. In our method, called Latent Unmixing, we apply predefined bandpass filters to the pixel latent space of the network output to pool this dimension into channels. This forces the network to learn a mapping that transforms the input overlapping distributions into a latent space where the distributions do not overlap and can be separated with bandpass filters (Fig. 1).
Algo. 2 describes the training process of the latent unmixing method. The training dataset must contain mixed volumetric data samples , along with unmixed ground truth images for each of ground truth channels. At each epoch, the latent representation of a first batch of data samples is generated with the 3D U-Net model (line 6). Predefined bandpass filters are applied to this latent vector to aggregate the output values included within each filter to the corresponding channel (lines 8 and 9). The resulting multi-channel prediction is normalized along the channel dimension so that each pixel is represented by its probability of belonging to this channel (line 11). The loss is the mean square error between the predicted () and ground truth () multi-channel maps of probabilities. The loss is backpropagated to update the weights of the network using stochastic gradient descent. This process is repeated over all batches of training data, and over a number of epochs .
4 Experiments and results
4.1 MNIST with simulated decay
We first apply the Latent Unmixing approach to untangle the signal from overlapping 4, 8, 9 digits and background in the 3D-MNIST dataset. In the images, the SNR (parameter in Algo. 1) is such that a simple 2D U-Net can not accurately separate the digits from the intensity input only (Supp. sec. D), demonstrating that the time dimension is needed to assign individual photons to the right channel. The decay parameters were chosen as 0.1, 0.2, 0.3 and 0.4 for digits 4, 8, 9 and the background respectively (Fig. 2, top-left). Our Latent unmixing method learns to map the input distributions sampled in the decay function to a pixel latent space where the contributions are well separated to fit the predefined bandpass filers (Fig. 2 and 3). The latent space is filtered with 4 bandpass filters of width 7 equally spaced along the pixel latent vector of length 28. The method can separate the digits accurately, achieving a Pearson correlation coefficient between true and predicted pixel intensities in the range above 0.82 for all channels over the 1,600 images from the test set (Fig. 4). To compare with curve fitting methods, we used MLE to unmix the same test set pixel-by-pixel. We fix the decay constants as the real values, and use MLE to fit the relative abundance. MLE achieved Pearson correlation coefficients between -0.1062 and 0.2484, showing no or negligible correlation between true and predicted intensities (Supp. sec. E and Table 2). This result was expected, since MLE is not an accurate method for separating as many as 4 colors with pixel counts in the range [0,20]; it is better suited for unmixing less colors with much higher pixel counts [3].
| Pearson corr. | MLE | Latent unmixing |
| channel 4 | -0.1062 | 0.8818 |
| channel 8 | 0.2484 | 0.8434 |
| channel 9 | 0.0426 | 0.8283 |
| background | 0.1331 | 0.9198 |
4.2 STED-FLIM
The simulated FLIM dataset serves as a demonstration of latent unmixing in a real experimental setting, in which channel separation is challenging for low photon counts or more than two fluorescent markers. The 3D U-Net model was trained to map the input volumes of depth time-bins to a latent space of the same length, on which 4 bandpass filters equally spaced and of equal width are applied, similarly to the 3D-MNIST experiment. Our latent unmixing approach accurately predicts the pixel-wise contributions of 4 components on the test images (Fig. 5) in conditions where MLE and phasor approaches fail at correctly separating each channel (Suppl. Fig. 6 and 7). The photon counts are too low (100) for MLE to fit exponential decay curves to individual pixels; it requires more than 100 photons per pixel to accurately separate 2-color images with known lifetimes of 2 and 5 nanoseconds (Supp. sec. F). The phasor method performs better than MLE, but does not reach the capacities of our method for 4-color unmixing. It is better suited for 2- or 3- color images, which our method can also learn to unmix by adapting the number and width of filters to the target number of channels (Supp. sec. F).
| MLE | Phasors | Latent unmixing | |
| 0.9031 | 0.5046 | 0.9716 | |
| 0.0023 | 0.2601 | 0.7635 | |
| 0.0115 | 0.2577 | 0.8855 | |
| -0.0568 | 0.1327 | 0.5539 |
4.2.1 Denoising with intermediate filters
One key particularity of the latent unmixing method is that the filters applied to the latent space can be displaced to fit the needs of the task and the distribution of the latent values. For 3D-MNIST, the output values of the pixel latent space tend to form distributions with a maximum aligned with the center of the filter for each channel (Fig. 2 and 3). In the STED-FLIM experiment, the distributions tend to be skewed towards the intersection of the filters (Fig. 6). This can be interpreted as the network not being able to correctly assign a channel to a given photon, placing its contribution at the intersection of two channel-filters. We tested this hypothesis by discarding at inference the photons that are assigned to the bins surrounding the intersection of each pair of channels. This creates a fifth channel composed solely of the photons that were not confidently classified in the latent space. By removing this channel, we get an unmixed and denoised version of the input volume (Fig. 6).
4.3 Mode decomposition for multi-mode optical fibers
The S-MMF dataset is used to test if the latent unmixing method can generalize to other types of distributions; here, additive sinusoidal signals with varying spatial intensities. For every single pixel, these signals are sampled under the Nyquist frequency, which is the minimal frequency of acquisition that would allow the application of Fourier-based processing methods. By applying our approach to the S-MMF volumes, the correlation from multiple pixels provide enough information to find the unknown intensity and spatial distribution of multiple coherent electromagnetic modes propagating simultaneously in a given optical fiber. The latent unmixing method was trained to map the input data to a latent space of length 60, where 6 bandpass filters of width 10 are applied to separate the space into 6 target modes (Fig. 7). Results demonstrate that the latent unmixing method can retrieve the spatial distributions of the modes (Fig. 8), which varies depending on the index profile of the fiber. To quantify the results, we compared the predicted relative intensity of each mode to the ground truth for a given sample volume. The total relative intensity is obtained by summing the pixel values over the predicted image of each mode. The predicted and ground truth fractions of intensity highly correlate 3 over all numbers of propagating modes. Interestingly, our method seems to perform better when the fiber has some intensity 4 or 5 modes than 2 or 3 modes. This goes against the trend of other methods that generally perform poorly for increased numbers of modes [24]. When there are more modes, the input volume includes more spectral variety and the 3D convolution is able to extract more information from cross-correlations gathered across the whole volume.
| Number of modes | Pearson |
| 1 mode | 0.9998 |
| 2 modes | 0.7551 |
| 3 modes | 0.8594 |
| 4 modes | 0.9567 |
| 5 modes | 0.9420 |
| 6 modes | 0.7598 |
5 Discussion and conclusion
We present latent unmixing, a novel approach for processing mixed images into their individual contributing components. Latent unmixing uses a 3D U-Net to combine the spatial and time dimensions in contexts where, if taken individually, each dimension (spatial and time) does not include enough information to allow correct separation of the components. Our approach takes advantage of an already demonstrated neural network architecture, the 3D U-Net, but the architecture choice is not a limiting factor of the method: it could easily be adapted to other architectures.
Our approach is first illustrated and validated on a toy synthetic dataset based on MNIST, where digit images are juxtaposed, generating the time dimension with different configurations for each of the class used. We then demonstrated the latent unmixing method for two experimental contexts: FLIM and mode decomposition in optical fibers, exemplifying the generalizability of the method to different distributions. In both cases, our method adapts to the fact that classes are unbalanced across data samples (number of photons per pixel in FLIM and number of propagating modes in MMF). In this study, we crafted datasets that accurately represent real experiments, while providing an associated ground truth to evaluate the performance of the proposed latent unmixing approach.
The latent unmixing method also offers interpretability opportunities that were briefly covered in this work. By mapping to a pixel latent space with more values than the number of channels in the ground truth (L=28 for four-channel MNIST and FLIM, L=60 for six-channel S-MMF), the network can learn to use the extra values to encode its certainty. A latent distribution that shifts from the middle of the filter can be interpreted as the network not being confident whether the photon belongs to one or the other channel. We studied this by adding narrow bandpass filters at the intersection between two filters, only at inference, for the FLIM dataset. Qualitative evaluation supports the hypothesis that photons leading to off-centered distributions seem to be associated with higher uncertainty from the network and can be used as a method of denoising images, as noise photons are more difficult to classify. This application is useful for real-life cases, like FLIM, where a background channel cannot be explicitly defined in the ground truth, as it could be for the MNIST dataset.
In conclusion, the latent unmixing method takes advantage of the capacity of deep multidimensional convolutional neural networks to combine the information from all dimensions, which allows unmixing data that is too under-sampled or noisy for each dimension to be processed independently. The method was crafted in a way that is generalizable to different number of target channels and different types of distributions. This is a novel use of deep neural networks, of great interesting for a range of imaging applications, showcasing the capacity of deep networks to process sparse information in its entirety.
Acknowledgments
We acknowledge the contributions of Andréanne Deschênes, Antoine Ollier and David Vincent for the scripts used to test the phasor method and maximum likelihood estimation, and Jean-Michel Bellavance for acquiring the STED images. This work was supported by the Neuronex Initiative (2014862) - National Science Foundation & Fond de Recherche du Québec - Santé, by the Natural Sciences and Engineering Research Council of Canada - Discovery Grant (DGECR-00200-2019), by the Canada Research Chair Tier 2 (CRC-0126-2019) and Canada-CIFAR AI Chair. CB and VB are both supported by NSERC PhD scholarships.
Supplementary Material
A. Acquisition of microscopy images
The STED-FLIM dataset was generated from real 4-colors fluorescence microscopy images acquired with a commercial Abberior STED system. Four proteins of interest were marked with fluorescent dyes that have excitation and emission spectral signatures that allow them to be acquired sequentially. The neuronal marker MAP2 was tagged with ATTO490LS, the postynaptic marker PSD95 with STAR635, the presynaptic marker VGLUT1 with Alexa488, and the RNA binding protein FUS with Alexa594. STAR635 and Alexa594 were imaged with the STED modality (resolution 20 nm), while ATTO490LS and Alexa488 were imaged with the confocal modality (resolution 200 nm).
B. Architecture and training parameters
Table 1 summarizes the architecture parameters for the 3D U-Net (Supp. fig. 1) used to obtain the results presented in this paper. All kernel sizes used thorughout the network are , and convolutions use zero-padding of size 1 along the 3 dimensions. Table 2 describes the layers inside the DoubleConv3D blocks.
| Block | Type | In Ch. | Out Ch. |
| 1 | DoubleConv3D | L | 64 |
| 2 | Down3D | 64 | 128 |
| 3 | Down3D | 128 | 256 |
| 4 | Down3D | 256 | 512 |
| 5 | Down3D | 512 | 1024 |
| 6 | Up3D | 1024 | 512 |
| 7 | Up3D | 512 | 256 |
| 8 | Up3D | 256 | 128 |
| 9 | Up3D | 128 | 64 |
| 10 | OutConv3D | 64 | 1 |
| Layer | Type | Channels |
| 1 | Convolution | In Ch. Out ch. |
| 2 | BatchNorm3d | Out Ch. Out ch. |
| 3 | ReLU | |
| 4 | Convolution | Out Ch. Out ch. |
| 5 | BatchNorm3d | Out Ch. Out ch. |
| 6 | ReLU |
Down3D blocks are a MaxPooling layer with kernels of shape and a stride of 2, followed by a DoubleConv3D block. Up3D blocks are DoubleConv3D blocks preceded by a 3D transposed convolution, also with kernels of shape and a stride of 2. The last convolution, OutConv3D, is a simple convolution with kernels of size and a stride of 1, to remove a dimension before applying the bandpass filters.
Table 3 summarizes the training hyper-parameters used for each dataset.
| 3D-MNIST | FLIM-STED | S-MMF | |
| Learning rate | 0.01 | 0.01 | 0.01 |
| Batch size | 64 | 8 | 16 |
| Input size | 2828 | 12828 | 6460 |
| Output size | 2828 | 12828 | 6460 |
| # Filters | 4 | 4 | 6 |
| Filter width | 7 | 7 | 10 |
| # images | 8,000 | 8,261 | 1,500 |
| Best epoch | 167 | 29 | 182 |
C. Generation of the 3D-MNIST dataset
The 3D-MNIST dataset was created by combining images of up to 3 digits from the MNIST dataset. To maximize the complexity of the unmixing process, we deliberately selected three digits characterized by the highest degree of overlap, ensuring that a substantial number of pixels would count multiple components. To find the trio of digits with the highest fraction of overlap, 1000 samples of digits 0, 1 and 2 were randomly sampled in the MNIST training set and the average intersection over union was computed. This process was repeated for all combinations of three digits, without repetitions of digits. The IOU for all triplets are shown in figure 2. The maximum average IOU (0.19) is obtained from the combinations of digits 4, 8 and 9.
D. 2D U-Net for unmixing MNIST
To make sure that the time dimension is exploited by the network to solve the unmixing task, a 2D version of the 3D-MNIST dataset was generated by omitting lines 23-26 and 30-33 in algo. 1; that is, by simply not creating the time dimension after creating the noisy overlapping digits. Results obtained by this 2D U-Net are shown in figure 3. The 2D U-Net can not separate the digits through the noise, demonstrating that the third dimension is useful for unmixing the overlapping digits.
E. MLE to unmix 3D-MNIST
Maximum likelihood estimation (MLE) is used as a baseline to demonstrate the difficulty of unmixing 4-color images with low photon counts and similar decay constants. MLE is used to fit the parameters A, B and C in equation 1 over the 3D-MNIST test set. The decay constants are fixed to help the method, as MLE is not suited for fitting as many as 7 parameters. The resulting confusion matrices are shown in fig. 4.
F. MLE and phasor method for STED-FLIM
Maximum likelihood estimation is used to fit the parameters A, B and C in equation 2 for the STED-FLIM test set.
The confusion matrices between true and predicted photon counts for each channel are shown in Supp. fig. 6(a). The MLE method tends to assign most photons to the shortest lifetime component, A.
We also applied MLE to fit the relative abundance of two components with lifetimes of 2 and 5 ns, which approximately correspond to the lowest and highest lifetime of dyes used for FLIM microscopy. We tested the method for pixel counts up to 10,000 to assess the number of photons required for MLE to correclty guess the relative abundance of both components, in the case where there are only two components in the pixel and their lifetimes are known (eq. 3). The predicted proportion for the first component (A) is shown in Supp. fig. 5 for varying numbers of photons and ground truth values of A.
The phasor analysis method was also used to unmix the 4-color FLIM-STED dataset, and the resulting confusion matrices are shown in Supp. fig. 6(b), with example results from the FLIM-STED test set in Supp. Fig. 8 and 9.
G. Code availability
All the code used to generate the 3D-MNIST dataset, train the models, and produce the results presented here is available at https://github.com/FLClab/LatentUnmixing. All experiments were conducted using Python 3.8.10, with the library versions indicated in the repository.

### 06 Apr 2024 v2

@@
-Filtering Pixel Latent Variables for Unmixing Volumetric Images
+
 Abstract
-Measurements of different overlapping components require robust unmixing algorithms to convert the raw multi-dimensional measurements to useful unmixed images.
-Such algorithms perform reliable separation of the components when the raw signal is fully resolved and contains enough information to fit curves on the raw distributions.
-In experimental physics, measurements are often noisy, undersampled, or unresolved spatially or spectrally.
-We propose a novel method where bandpass filters are applied to the latent space of a multi-dimensional convolutional neural network to separate the overlapping signal components and extract each of their relative contributions.
-Simultaneously processing all dimensions with multi-dimensional convolution kernels empowers the network to combine the information from adjacent pixels and time- or spectral-bins, facilitating component separation in instances where individual pixels lack well-resolved information.
-We demonstrate the applicability of the method to real experimental physics problems using fluorescence lifetime microscopy and mode decomposition in optical fibers as test cases.
-The successful application of our approach to these two distinct experimental cases, characterized by different measured distributions, highlights the versatility of our approach in addressing a wide array of imaging tasks.
+The development of robust signal unmixing algorithms is essential for leveraging multimodal datasets acquired through a wide array of scientific imaging technologies, including hyperspectral or time-resolved acquisitions.
+In experimental physics, enhancing the spatio-temporal resolution or expanding the number of detection channels often leads to diminished sampling rate and signal-to-noise ratio, significantly affecting the efficacy of signal unmixing algorithms.
+We propose applying band-pass filters to the latent space of a multi-dimensional convolutional neural network to disentangle overlapping signal components, enabling the isolation and quantification of their individual contributions.
+Using multi-dimensional convolution kernels to process all dimensions simultaneously enhances the network’s ability to extract information from adjacent pixels, time- or spectral-bins.
+This approach enables more effective separation of components in cases where individual pixels do not provide clear, well-resolved information.
+We showcase the method’s practical use in experimental physics through two test cases that highlight the versatility of our approach: fluorescence lifetime microscopy and mode decomposition in optical fibers.
+The latent unmixing method extracts valuable information from complex signals that cannot be resolved by standard methods.
+Application of latent unmixing to real FLIM experiments will increase the number of distinguishable fluorescent markers.
+It will also open new possibilities in optics and photonics for multichannel separations at increased sampling rate.
 1 Introduction
-Unmixing refers to the technique of breaking down a measurement into its unique constituents and their respective proportions.
-This concept is commonly linked to spectral analysis, where the overlapping spectral signatures of different emitters hinder the accurate identification of each emitter’s abundance, especially when the spectral signature is undetermined or the data is insufficiently sampled.
-Beyond spectral measurements, unmixing applies to any measurements that combine spatial dimensions (x,y,z) to additional continuous and often multi-channel measures.
-In fluorescence lifetime microscopy, this additional dimension is the time it took for the photon to be emitted[4].
-In spectrally and spatially (S) resolved measurements for characterizing optical fibers, this dimension is the amplitude of the propagated light for different input wavelengths [21].
-In this work, we propose a generalized unmixing method based on convolutional neural networks that applies to various types of multi-dimensional measurements (Fig.
+Signal unmixing refers to the technique of breaking down a measurement into its unique constituents and their respective proportions.
+This concept is often associated with spectral analysis, where the challenge lies in accurately identifying the abundance of each emitter from their overlapping spectral signatures.
+Beyond spectral data, signal unmixing can apply to any measurements that rely on the combination of multiple dimensions or components.
+The difficulty to accurately separate the components becomes more pronounced when the specific spectrum of an emitter is unknown or when the data coverage is insufficient.
+In this work, we propose an unmixing method that applies to various types of underdefined multi-dimensional measurements that outperforms classical curve-fitting based methods such as Maximum Likelihood Estimation[4].
+Our method, called Latent Unmixing, transforms the overlapping input contributions to a latent space where they are untangled and can be separated by applying filters directly in the latent space (Fig.
 1).
-We specifically focus on noisy data or undersampled distributions, for which common methods fail to accurately unmix the different contributions.
-The proposed method relies on transforming the overlapping input distributions to a latent space where they are untangled and can be separated using simple bandpass filters.
-We demonstrate the method on synthetic datasets based on real data and created for real experimental physics use cases, showcasing the general applicability of the approach.
-2 Related work
-In this work, we demonstrate an unmixing method on two experimental physics use cases: fluorescence lifetime microscopy (FLIM) and mode decomposition in optical fibers.
-In both cases, the input data is composed of overlapping distributions that need to be decomposed into single channel outputs to retrieve useful information.
-2.1 Multidimensional CNNs and the latent space
-Three-dimensional Convolutional Neural Networks (3D CNNs) offer advantages over their two-dimensional counterparts in tasks involving volumetric imaging such as medical imaging [28] or spatiotemporal data such as videos [14].
-Their ability to capture both spatial and temporal features simultaneously leads to improved representations of complex patterns in three-dimensional space [32].
-Unmixing tasks face the challenge of combining 3D input with 2D multi-channels output, making the task difficult to adapt to existing 3D models; for example, with a 3D U-Net, where the 3D features learned in the encoder cannot be directly linked to the 2D features required in the decoder.
-CNNs that take as input a 3D volume and predict a 2D image is not a common architecture as it is a distinct task from volumetric segmentation, and no architecture has been developed specifically for this task.
+1.1 Related work
+Linear unmixing methods exploit the fact that the signal measured at each pixel is a linear combination of each components’ specific signature weighted by their relative abundance.
+When the signature is known, the relative abundance parameters can be retrieved using estimation algorithms such as least squares fitting methods[41, 19] and Maximum Likelihood Estimation (MLE) [4]; given that the signal is sufficiently sampled for the curves to be accurately estimated from the raw data.
+Another linear unmixing approach, the Phasor method[24], translates the raw measurements to a 2D phasor space where the distributions are represented by the real and imaginary components of the Fourier transform.
+This representation then allows to estimate the relative abundance of different components through a linear interpolation, knowing the phasor coordinates of individual components obtained through control measurements.
+Deep learning methods have been proposed for the task of unmixing images, especially for hyperspectral data[17, 8, 50].
+Hyperspectral data has a high spectral sampling obtained at the expense of diminished spatial resolution[46].
+Although this limitation poses its own challenges, mainly that single pixels cover large ground areas and therefore often contain several materials [51, 14], the methods developed do not cover the inverse problem of measurements that are well resolved spatially but undersampled spectrally.
+The deep learning models for processing hyperspectral data – either for classification[1, 53, 23], super-resolution[26], denoising[40], or unmixing[47, 34] – are generally based on three-dimensional Convolutional Neural Networks (3D CNNs) to process the raw hyperspectral measurements as volumetric data.
+3D CNNs offer advantages over their two-dimensional counterparts in tasks involving volumetric datasets [45] or spatiotemporal acquisitions [22].
+Their ability to capture and combine spatial and temporal features simultaneously leads to improved representations of complex patterns in three-dimensional space [49].
+For signal unmixing of 3D imaging data, it is required that the CNN takes as input a 3D volume and predicts the proportion of n component in each pixel, resulting in a stack of n 2D images.
+Signal unmixing tasks present the significantly different challenge of converting a 3D input (e.g.
+(x, y, t)) into a 2D multi-channel output (e.g.
+n (x, y)), which can not directly be solved with existing end-to-end 3D models, such as the 3D U-Net.
 Tasks on 3D volumes are generally either in the volumetric space from input to output (3D segmentation [45]) or a classification layer is added to the learned representation for a 1D output (classification [33]).
-2.5D networks have been proposed for the segmentation of volumes into 2D slice-wise segmentation for cases where 3D CNNs are not applicable because of slice discontinuities [23], or in cases where the data is too limited to train 3D networks [15].
-However, 2.5D networks were shown to be less accurate than 3D CNNs when 3D CNNs are applicable [2].
-In our method, the features are represented in the 3D space from the entry to the exit of the network, so that existing 3D architectures can be exploited for unmixing tasks.
-This leads to a general model that is not constrained in the number of target output channels, because the learned features are translated into channels directly from the learned 3D latent space.
-This latent space is an abstract and simplified representation of the data; in convolutional neural networks, the latent space simplifies the raw input pixel data into an informative representation that is used by the subsequent layers to perform the task[17].
-2.2 Fluorescence lifetime microscopy
-Fluorescence lifetime microscopy (FLIM) uses the arrival time of fluorescence photons to assign them to specific fluorophores or to measure environmental changes in biological samples [4].
-The fluorescence decay follows an exponential probability distribution characterized by the parameter corresponding to the fluorescence lifetime [5].
-The fluorescence lifetime represents a distinctive photophysical characteristic unique to each fluorescent molecule, which is used in FLIM to discriminate between several fluorophores.
-However, even by carefully selecting molecules based on their fluorescence lifetime, exponential distributions always overlap and signal unmixing can be challenging, especially in the low photon-count regime (100 photons/pixel) or when multiple fluorophores are present at a given position.
-For live-cell imaging, where photobleaching and sample movement should be minimized, FLIM measurements are generally associated with low photon-counts [31].
-Additionally, measurements of interactions between multiple components increase the complexity of assigning each photon to a given emitting molecule.
-2.2.1 FLIM analysis
-Different unmixing approaches have been proposed for FLIM such as Maximum Likelihood Estimation (MLE) [3], phasor analysis [8], and deep learning based strategies [30].
-MLE is an effective method for fitting parameters to describe a curve given a set of samples from the distribution and is a common approach for FLIM data analysis [3].
-An alternative approach relying on phasor analysis has been proposed for unmixing fluorescence lifetime distributions [8].
-The performance of both approaches is strongly limited by low photon counts (300) and when more than three components have to be unmixed [7, 6, 25, 8].
-Recently, deep learning based approaches have been proposed for FLIM data analysis.
-The Fluorescence lifetime imaging network (FLI-Net) [30] is a 3D-CNN that takes as input a 3D FLIM volume (x,y,t) where each pixel is characterized by photons sampled in a biexponential distribution with aim to assign a mean lifetime to each pixel rather than unmixing the input image.
-The Fluorescence Lifetime AdderNet is a 1D network trained for lifetime unmixing [33], which can handle spatially distinct structures and does not use the spatial features to improve photon assignment in overlapping structures.
-2.2.2 Unmixing strategies for fluorescence microscopy
-A common method for unmixing multi-channel images from a single acquisition is to combine the lifetime and spectral information (spectral FLIM, S-FLIM).
-In this case, acquired volumes are 4D (x,y,,t).
-The multidimensional information facilitates the assignment of photons to a given fluorescent molecule, since emitters characterized by similar lifetimes can have different spectral signatures or vice versa.
-Phasor analysis has been explored as a method for unmixing S-FLIM data [19, 10].
-Deep learning-based unmixing of S-FLIM data has also been demonstrated with UNMIX-ME [29], which uses convolutional neural networks to map noisy hyperspectral FLIM data to spatially independent unmixed fluorescence coefficient values.
-The method allows the separation of up to four components in simulated samples.
-To our knowledge, there is a gap in the literature for FLIM unmixing method that applies to images in the low photon count regime (100 photon/pixel) and more than two fluorescent emitters per pixel.
-Such an approach would be transformative for the democratization of FLIM for multi-channel live-cell fluorescence microscopy and nanoscopy.
-2.3 Mode decomposition in multi-mode fibers
-Multi-mode fibers are optical fibers that support the propagation of multiple simultaneous light modes [11].
-Spatially and spectrally resolved (S) imaging[21] is a measurement commonly used for characterizing how light travels through multi-mode optical fibers.
-By measuring the spatial distribution of the intensity at the output of the fiber for a range of laser wavelengths, the modes of propagation of the light through the fiber can be characterized.
-The 2D image at the fiber output is not merely a linear combination of each distinct mode because the coherent modes interfere constructively or destructively with each other in function of their relative phases.
-In a two modes fiber, the intensity of a single pixel versus the wavelength of the laser propagating through the fiber, within a small spectral range, oscillates periodically with a sinusoidal distribution (Fig.
-7).
-When more than one mode propagates through the fiber, the intensity oscillates as the sum of each sinus function associated to each combination of two modes.
-In a properly sampled measurement where the frequency of acquisition is above the Nyquist frequency, these oscillations allow to separate each mode with a Fourier transform based method [20] or by principal component analysis (PCA) [24, 26].
-Other methods exist to achieve real-time mode decomposition from a single image with various algorithms [13, 1, 18], but they all require the a priori knowledge of the spatial distribution of each mode.
-To our knowledge, no method for mode decomposition in multi-mode fibers solve both limitations of current approaches: separate modes with unknown spatial distributions, and successfully retrieve the contribution of each mode with a sub-Nyquist sampling rate.
+We designed our method over the intuition that a 3D U-Net can be optimized to map the input data from the input space, where the signal components are undersampled and overlapping, to a space where the different contributions are distinctly differentiated.
+Other works previously used the distangled representation of input data in the latent space of an autoencoder to unmix hyperspectral data[38, 32].
+Those methods are however specific to hyperspectral data, because they either require precise reference spectra[38] and therefore do not account for possible spectral variability[9], or process a single pixel at a time and do no use the spatial and spectral information from neighbouring pixels[32].
+For hyperspectral data, the pixel spatial range is sufficiently large that there may be no need to rely on information from neighboring pixels to infer the components present within the pixel.
+1.2 Contributions
+The differentiating contributions of this work are as follows: 1) we developed a method to map input volumetric measurements to a latent space where the included signals are disentangled, 2) we specifically focus on the less studied unmixing problem for data that is well-resolved spatially but undersampled in its third dimension and outperform commonly used methods for low-sampling regimes, 3) we showcase the applicability of the method over three simulated datasets for two completely distinct experimental use cases and include a demonstration on real acquired data, which is enabled by the generalizability of the model to different parametrization of the input distributions.
 2 Methodology
-3.1 Datasets
-Experiments are conducted with three datasets: 1) a modified version of the MNIST dataset [16] with overlaid digits and simulated lifetime (3D-MNIST), 2) a FLIM dataset generated from STimulated Emission Depletion (STED) microscopy images of neuronal proteins (FLIM-STED), and 3) a simulation of spectrally and spatially resolved measurements of light propagating in multi-mode fibers (S-MMF).
-3D-MNIST The 3D-MNIST dataset was used to develop the method and evaluate the influence of different parameters.
-Images from the classes 4, 8 and 9 are overlaid to create mixed images.
-Random Poisson noise is generated over the whole image as a fourth target channel.
-This multi-dimensional version of MNIST is generated according to Algo.
-1.
-The generated MNIST dataset is parametrized by a signal-to-noise factor , decay constants for each channel (, , , ), a time dimension length , and a number of images .
-Each data sample is created by randomly sampling up to three digits from classes (4, 8, 9) from the MNIST training set.
-We chose the digits 4, 8 and 9 because this triplet of digits has the highest fraction of overlap (Supp.
-fig.
-2).
-There is a probability 0.5 of picking one digit for each of the (4,8,9) classes (lines 9 to 13), such that a training sample can contain 0, 1, 2 or 3 digits.
-The images are independently randomly rotated, flipped and translated (line 17).
-To reduce the pixel counts from [0,255] to a noisier range more representative of real experimental conditions, the image is first normalized to the range [0,1] by dividing each image by its maximum value (line 19).
-A random count is assigned to each pixel by sampling an intensity value in a Poisson distribution parameterized by the pixel value, multiplied by the SNR factor (lines 21 and 22).
-For each count, an equivalent number of time values are sampled following an exponential distribution parameterized by the decay constant of the channel (line 24).
-The time values of each individual pixel are converted to a histogram of bins of fixed width (line 26).
-A background channel is also created by randomly sampling values in a Poisson distribution with parameter and creating a time histogram following the same method as for the digits channels (lines 29 to 33).
-The four volumes (4, 8, 9, background) are added along the channel dimension to create a mixed volume of dimensions () (line 35).
-The process is repeated times to generate the 3D-MNIST dataset.
-The 3D-MNIST dataset counts 8,000 combined images generated with and , from which 6,400 randomly sampled images were used for training and 1,600 for testing.
-STED-FLIM The STED-FLIM dataset consists of real microscopy images of four different proteins tagged with fluorescent markers in fixed cortical neurons (Suppl.
-sec.
-A).
-The signal from each fluorescent marker was spectrally separated to obtain ground truth images with four independent channels.
-The STED-FLIM dataset was generated by adapting the 3D-MNIST data generation algorithm (algo.
-1).
-The following modifications were applied to the algorithm: 1) dataset is a set of 4-color FLIM images, 2) channels are not randomly discarded and independently sampled (lines 10-15), 3) the true measured photon counts were preserved instead of normalized and randomized (lines 16-22), 4) no background channel was added (lines 29-33).
-The lifetime values (Suppl.
-table 1) were chosen to fit within the range of commercially available fluorescent dyes for STED microscopy [12].
-We chose values within this range so that the separation between lifetimes is variable (0.7 ns A-B, 0.6 ns B-C, 0.2 ns C-D).
-Since the lifetime distributions are synthetic, the actual lifetimes do not need to match a specific dye, so we attributed generic names (A, B, C and D) to the fluorescent markers for clarity.
-Channels A and C were acquired with confocal microscopy while channels B and D were acquired with STED microscopy.
-This allows to test if the method can adapt to varying photon counts and spatial resolutions.
-The STED-FLIM dataset counts 400 training and validation images which were preprocessed into 8,261 crops, and 20 test images.
-| Protein | Marker | Lifetime [ns] |
-| MAP2 | A | 2.6 |
-| PSD95 | B | 3.3 |
-| VGLUT1 | C | 3.9 |
-| FUS | D | 4.1 |
-S-MMF The S-MMF dataset consists in images of the intensity measured at the output of optical fibers for different input wavelengths [22].
-The simulated measure has three dimensions; two spatial (2D images of the light distribution at the fiber output) and one spectral (each image represents the measure obtained from a different light wavelength).
-The dataset was computed with the equations described in Quiquempois et al.
-(2021) [24].
-A vector finite difference mode-solver [9] was used to find mode properties around 1550 nm for various index profiles and fiber geometries.
-All fibers in the dataset support up to 6 modes.
-The mode spatial distribution and effective index dependency on wavelength (dispersion) were taken into account.
-For all data samples, we considered a spectral range of 6 nm for a fiber length of 5 m.
-The dataset includes measurements with various combinations of up to 6 propagating modes, four different fiber geometries (step-index with varying numerical apertures and core size, GRIN, W-type, H-type and a realistic step-index geometry with a central dip), and different mode-specific intensities ().
-We used 1500 fiber measurements to train the network and 175 for testing.
 2.1 Model
-Network architecture 3D convolution kernels allow the processing of all dimensions of a 3D input volume simultaneously.
-We chose to use a 3D CNN for unmixing undersampled data so that the 3D kernels can combine essential information from neighboring pixels and neighboring time- or spectral-bins.
-The network used here is a standard 3D U-Net (Supp.
+We chose to use a 3D U-Net for unmixing undersampled data so that the 3D kernels can combine essential information from neighboring pixels and neighboring time- or spectral-bins, since 3D convolution kernels allow the processing of all dimensions of a 3D input volume simultaneously.
+The network used here is a 3D U-Net with 4 downsampling blocks and 4 upsampling blocks (Suppl.
 sec.
 B).
+To generalize the model to contexts where the input distribution parameters vary, we developed a conditional paradigm where the parameters that define the distribution of the input volume is embedded to the encoder blocks.
+The whole architecture is described in details in Supp.
+sec.
+B.
 As is standard for U-Nets, the dimensionality of the input remains unchanged through the layers; the input 3D volume results in a 3D output of the same size.
 We added a layer with one convolution filter at the end of the network that maps the output of the U-Net to a single-channel volumetric output of dimension , with and coming from the images spatial dimensions and being the pixel latent space size.
-Latent space filters For unmixing applications, the input is a 3D volume and the target is a channel 2D image.
+The receptive field of the final layer spans 110 pixels in width, height, and depth, contrasting with pixelwise approaches for which the predictive capacity is restricted by the very limited information of a single pixel.
+2.2 Latent space filters
+In convolutional neural networks, the latent space simplifies the raw input pixel data into an informative representation that is used by the subsequent layers to perform the task[25].
+Our latent unmixing method uses this abstract and simplified representation of the data in the latent space to translate the 3D input into a multi-channel output.
 In our method, called Latent Unmixing, we apply predefined bandpass filters to the pixel latent space of the network output to pool this dimension into channels.
-This forces the network to learn a mapping that transforms the input overlapping distributions into a latent space where the distributions do not overlap and can be separated with bandpass filters (Fig.
+This forces the network to learn a mapping that transforms the input overlapping distributions into a latent space where they do not overlap and can be separated with bandpass filters (Fig.
 1).
@@
 This process is repeated over all batches of training data, and over a number of epochs .
-4 Experiments and results
-4.1 MNIST with simulated decay
+3 Experiments
+In this work, we demonstrate a signal unmixing method on a variation of the MNIST dataset and on two experimental physics use cases: fluorescence lifetime microscopy (FLIM) and mode decomposition in optical fibers.
+In all three cases, the input data is composed of noisy, undersampled, and overlapping distributions that need to be decomposed into multi-channel outputs.
+3.1 3D MNIST with simulated decay
+Dataset The 3D-MNIST dataset was used to benchmark the method’s performance.
+Images from the digit classes 4, 8 and 9 are overlaid to create mixed images.
+We chose the digits 4, 8 and 9 because this triplet of digits has the highest fraction of overlap (Supp.
+fig.
+2).
+Random Poisson noise is generated over the whole image as a fourth target channel to simulate background noise in optical microscopy.
+A multi-dimensional version of the mixed images of the MNIST digits is generated according to Algo.
+2.
+The generated dataset is parametrized by a signal-to-noise factor , a decay constants for each channel (, , , ), a time dimension histogram of bins, and a number of images .
+Each mixed image is created by randomly sampling up to three digits from classes (4, 8, 9) from the original MNIST training set.
+There is a probability 0.5 of picking one digit for each of the (4,8,9) classes (lines 9 to 13), such that a training sample can contain 0, 1, 2 or 3 digits.
+The digits are independently randomly rotated, flipped and translated (line 17).
+The mixed images are first normalized to the range [0,1] by dividing each image by its maximum value (line 19).
+To more accurately mirror real experimental conditions, normalized pixel values are adjusted by sampling a random count for each pixel from a Poisson distribution, with the sampling parameter being the pixel value multiplied by the signal-to-noise ratio (SNR) factor (lines 21 and 22).
+For each count, a corresponding time is sampled following an exponential distribution parameterized by the decay constant of the channel (line 24).
+The time values of each individual pixel are converted to a histogram of bins of fixed width (line 26).
+A background channel is also created by randomly sampling values in a Poisson distribution with parameter and creating a time histogram following the same method as for the digits channels (lines 29 to 33).
+The four volumes (4, 8, 9, background) are added along the channel dimension to create a mixed volume of dimensions () (line 35).
+The process is repeated times to generate the 3D-MNIST dataset.
+The 3D-MNIST dataset counts 8,000 mixed volumes generated with and , from which 6,400 randomly sampled images were used for training and 1,600 for testing.
+3.1.1 Results
 We first apply the Latent Unmixing approach to untangle the signal from overlapping 4, 8, 9 digits and background in the 3D-MNIST dataset.
 In the images, the SNR (parameter in Algo.
-1) is such that a simple 2D U-Net can not accurately separate the digits from the intensity input only (Supp.
+2) is such that a simple 2D U-Net can not accurately separate the digits from the intensity input only (Supp.
 sec.
@@
 2, top-left).
-Our Latent unmixing method learns to map the input distributions sampled in the decay function to a pixel latent space where the contributions are well separated to fit the predefined bandpass filers (Fig.
-2 and 3).
-The latent space is filtered with 4 bandpass filters of width 7 equally spaced along the pixel latent vector of length 28.
-The method can separate the digits accurately, achieving a Pearson correlation coefficient between true and predicted pixel intensities in the range above 0.82 for all channels over the 1,600 images from the test set (Fig.
-4).
-To compare with curve fitting methods, we used MLE to unmix the same test set pixel-by-pixel.
-We fix the decay constants as the real values, and use MLE to fit the relative abundance.
+Our Latent Unmixing method learns to map the input distributions sampled in the decay function to a pixel latent space where the contributions are well separated to fit the predefined bandpass filters (Fig.
+2).
+Four bandpass filters of width 7 equally spaced along the pixel latent vector of length 28 are used.
+The method can separate the digits accurately, achieving a Pearson correlation coefficient between true and predicted pixel intensities in above 0.82 for all channels over the test set (Fig.
+LABEL:fig:mnistb).
+We compared the result of Latent Unmixing with the established curve fitting approach MLE to unmix the same test set pixel-by-pixel.
+We fix the decay constants as the actual values, and use MLE to fit the relative abundance.
 MLE achieved Pearson correlation coefficients between -0.1062 and 0.2484, showing no or negligible correlation between true and predicted intensities (Supp.
@@
 E and Table 1).
-This result was expected, since MLE is not an accurate method for separating as many as 4 colors with pixel counts in the range [0,20]; it is better suited for unmixing less colors with much higher pixel counts [3].
+This result was expected, since MLE is not an accurate method for separating as many as 4 channels with limited pixel counts in the range [0,20] [4].
 | Pearson corr. | MLE | Latent Unmixing |
@@
 | background | 0.1331 | 0.9198 |
-4.2 STED-FLIM
-The simulated FLIM dataset serves as a demonstration of latent unmixing in a real experimental setting, in which channel separation is challenging for low photon counts or more than two fluorescent markers.
-The 3D U-Net model was trained to map the input volumes of depth time-bins to a latent space of the same length, on which 4 bandpass filters equally spaced and of equal width are applied, similarly to the 3D-MNIST experiment.
+3.2 Fluorescence lifetime microscopy
+Fluorescence lifetime microscopy (FLIM) uses the arrival time of fluorescence photons to assign them to specific fluorophores or to measure environmental changes in biological samples [6].
+The fluorescence decay follows an exponential probability distribution characterized by the parameter corresponding to the fluorescence lifetime [7].
+The fluorescence lifetime is a unique photophysical property of each fluorescent molecule, utilized in FLIM to differentiate between fluorophores.
+Despite careful selection of fluorophores with distinct fluorescence lifetime, overlapping exponential distributions pose challenges for signal unmixing.
+This is particularly difficult with low photon counts (less than 100 photons/pixel) or when multiple fluorophores coexist at the same location[43, 28].
+For live-cell imaging, where photobleaching and sample movement should be minimized, FLIM measurements are generally associated with low photon-counts (¡25 photons per pixel[2]).
+3.2.1 Related work
+Different unmixing approaches have been proposed specifically for FLIM such as linear phasor analysis [13], and deep learning based strategies [42, 11, 54].
+Maximum Likelihood Estimation (MLE) is a well-established curve fitting approach, that requires known components distributions when applied to multi-channel unmixing of FLIM data [4].
+Phasor analysis has been proposed for the unmixing of fluorescence lifetimes, utilizing graphical representations to simplify the identification of composite lifetimes without requiring iterative fitting [13, 36].
+The performance of both approaches is strongly limited by low photon counts (300) and when more than three components have to be unmixed [12, 10, 37, 13].
+Recently, deep learning based approaches have been proposed for FLIM data analysis.
+The Fluorescence lifetime imaging network (FLI-Net) [42] is a 3D-CNN that takes as input a 3D FLIM volume (x,y,t) where each pixel is characterized by photons sampled in an exponential distribution with aim to assign a mean lifetime to each pixel rather than unmixing the input image.
+The fluorescence lifetime imaging method based on Generative Adversarial Network Estimation (FlimGANE)[11] uses a GAN to generate high-quality FLIM images based on photon-starved (50-200 photons per pixel) measurements.
+Both FLI-Net and FlimGANE models are limited to two components unmixing, since their architecture is designed for three outputs: lifetime values and and relative abundance .
+The lifetime of a fluorescent dye is not a fixed value and is known to vary based on the environmental conditions of the sample[44].
+Training a model on a predefined set of lifetime values is impracticable, as it would unlikely capture the true measured distributions of fluorescence lifetime in a real sample.
+In recent years, major developments have been made to enable multiplexed fluorescence microscopy, enabling the characterization of cellular and molecular interactions between multiple partners [16, 48, 5].
+Development of analytical methods that can be applied to limited photon budget in multiplexed FLIM imaging would have an important impact in the field, allowing to increase sampling rates for live-cell imaging or to resolve nanostructures in super-resolution microscopy.
+Our proposed approach addresses the challenge of unmixing over up to four independent fluorophores in FLIM data, tailored for contexts where the photon count per pixel aligns with that observed in STimulated Emission Depletion (STED) super-resolution microscopy or time-lapse confocal microscopy.
+3.2.2 Dataset
+The STED-FLIM dataset consists of 30 real microscopy images of four different proteins tagged with fluorescent markers in fixed cortical neurons (Suppl.
+sec.
+A).
+The signal from each fluorescent marker was spectrally separated to obtain ground truth images with four independent channels.
+Prior to training, the images were split into 256 256 pixels crops using a sliding window with steps of 128 pixels.
+Crops were kept in the training set if the mean photon count of each of the 4 channels was at least 1 photon per pixel.
+This operation resulted in a set of 1806 crops for training.
+To augment the data, channels were randomly permuted, and up to two channels were removed.
+The time dimension was created during training by first uniformly sampling lifetime values , , , , and then sampling random time values for each pixels inside the dataloader from an exponential decay curve parameterized by the channels’ lifetime.
+This realistic dataset format could easily be adapted to fit specific parameters, number of channels, and ranges of lifetime values.
+3.2.3 Training
+The STED-FLIM dataset serves as a demonstration of Latent Unmixing in a real experimental setting, in which channel separation is challenging for low photon counts (¡ 50 photons per pixel), for more than two fluorescent markers, and for small lifetime difference between the emitters (¡ 1 ns).
+To improve the generalizability of the model, we trained a conditional U-Net inspired by the denoising diffusion model[20] that takes as input a vector in addition to the volumetric image to unmix.
+For targets with less than 4 channels, the empty channels are represented by .
+is projected to the encoding blocks by passing through two learnable fully-connected layers both with 64 nodes with a GELU activation in-between and a SiLU activation afterwards.
+The learned representation then goes through a third fully-connected layer with an output size corresponding to the number of channels of each encoding block (128, 256, 512 and 1024).
+The resulting values are then used to scale and shift the input of each encoding block.
+We used the implementation from https://github.com/lucidrains/denoising-diffusion-pytorch/.
+This model was trained on the FLIM-STED dataset with varying lifetime values and resulted in a more general model that can be applied to a wider range of lifetime pairs than an unconditional model (Suppl.
+sec.
+G).
+The conditional 3D U-Net model was trained to map the input volumes of depth time-bins to a latent space of the same length, on which 4 bandpass filters equally spaced and of equal width are applied, similarly to the 3D-MNIST experiment.
+3.2.4 Results
+Robustness to low-photon counts
+(Fig. 4)
+Robustness to close lifetime values
+Finding pairs of fluorophores that are compatible with the proteins to be labeled and have distant lifetime values can be a challenge.
+The ability to correctly unmix two channels with lifetime that are close apart (¡1 ns) is therefore a valuable advantage for an unmixing method, since it opens up more possibilities of dyes that can be paired up.
+Our latent unmixing method shows better robustness to close lifetime than the phasor analysis method (Fig.
+5), achieving a Pearson correlation of for lifetime values only 0.1 ns apart, compared to the phasor analysis method that achieves only .
+4-color unmixing
 Our Latent Unmixing approach accurately predicts the pixel-wise contributions of 4 components on the test images (Fig.
-5) in conditions where MLE and phasor approaches fail at correctly separating each channel (Suppl.
-Fig.
-6 and 7).
-The photon counts are too low (100) for MLE to fit exponential decay curves to individual pixels; it requires more than 100 photons per pixel to accurately separate 2-color images with known lifetimes of 2 and 5 nanoseconds (Supp.
+6) in conditions where MLE and phasor approaches fail at correctly separating each channel (Suppl.
 Sec.
-F).
-The phasor method performs better than MLE, but does not reach the capacities of our method for 4-color unmixing.
-It is better suited for 2- or 3- color images, which our method can also learn to unmix by adapting the number and width of filters to the target number of channels (Supp.
+F.
+MLE and phasor method for STED-FLIM).
+3.2.5 Application to real FLIM measurements
+A key feature of the Latent Unmixing method is that it is trained on simulated data, therefore erasing the need of acquiring a FLIM dataset specifically for training.
+The simulated data can cover any range and combinations of lifetime values, enabling the training of a general model that can be applied to any pair of lifetime values.
+We applied the Latent Unmixing model trained on the simulated FLIM-STED dataset on real FLIM data (Fig.
+7) with different lifetime values, photon counts, biological structures, and number of channels to separate (2, 3 and 4); all using the same general model (Suppl.
 sec.
-F).
-| MLE | Phasors | Latent unmixing | |
-| 0.9031 | 0.5046 | 0.9716 | |
-| 0.0023 | 0.2601 | 0.7635 | |
-| 0.0115 | 0.2577 | 0.8855 | |
-| -0.0568 | 0.1327 | 0.5539 |
-4.2.1 Denoising with intermediate filters
-One key particularity of the latent unmixing method is that the filters applied to the latent space can be displaced to fit the needs of the task and the distribution of the latent values.
-For 3D-MNIST, the output values of the pixel latent space tend to form distributions with a maximum aligned with the center of the filter for each channel (Fig.
-2 and 3).
-In the STED-FLIM experiment, the distributions tend to be skewed towards the intersection of the filters (Fig.
-6).
-This can be interpreted as the network not being able to correctly assign a channel to a given photon, placing its contribution at the intersection of two channel-filters.
-We tested this hypothesis by discarding at inference the photons that are assigned to the bins surrounding the intersection of each pair of channels.
-This creates a fifth channel composed solely of the photons that were not confidently classified in the latent space.
-By removing this channel, we get an unmixed and denoised version of the input volume (Fig.
-6).
-4.3 Mode decomposition for multi-mode optical fibers
+LABEL:).
+The real measurements require only very simple preprocessing steps to fit the format of the simulated training data (Suppl.
+sec.
+LABEL:).
+3.3 Mode decomposition in multi-mode fibers
+Multi-mode fibers are optical fibers that support the propagation of multiple simultaneous light modes [18].
+Spatially and spectrally resolved (S) imaging[31] is a measurement commonly used for characterizing how light travels through multi-mode optical fibers.
+By measuring the spatial distribution of the intensity at the output of the fiber for a range of laser wavelengths, the modes of propagation of the light through the fiber can be characterized.
+The 2D image at the fiber output is not merely a linear combination of each distinct mode because the coherent modes interfere constructively or destructively with each other in function of their relative phases.
+In a two modes fiber, the intensity of a single pixel versus the wavelength of the laser propagating through the fiber, within a small spectral range, oscillates periodically with a sinusoidal distribution (Fig.
+8).
+When multiple modes propagate through the fiber, the intensity oscillates as a composite of sinusoidal functions, each corresponding to the interaction between pairs of modes.
+In an measurement for which the sampling frequency is above the Nyquist criterion, the measured oscillations allow to separate each mode with a Fourier transform based method [30] or by principal component analysis (PCA) [35, 39].
+Other methods exist to achieve real-time mode decomposition from a single image with various algorithms [21, 3, 27], but they all require the a priori knowledge of the spatial distribution of each mode.
+To our knowledge, no method for mode decomposition in multi-mode fibers allows to separate modes with unknown spatial distributions and successfully retrieve their contribution in measurements using a sub-Nyquist sampling rate.
+3.3.1 Dataset
+The S-MMF dataset consists in simulated images of the intensity measured at the output of optical fibers for different input wavelengths [NichoModedecompositionformulti-modeopticalfiberslson:08].
+The simulated measure has three dimensions; two spatial (2D images of the light distribution at the fiber output) and one spectral (each image represents the measure obtained from a different light wavelength).
+The dataset was computed with the equations described in Quiquempois et al.
+(2021) [35].
+A vector finite difference mode-solver [15] was used to find mode properties around 1550 nm for various index profiles and fiber geometries.
+All fibers in the dataset support up to 6 modes.
+The mode spatial distribution and effective index dependency on wavelength (dispersion) were taken into account.
+For all data samples, we considered a spectral range of 6 nm for a fiber length of 5 m.
+The dataset includes measurements with various combinations of up to 6 propagating modes, four different fiber geometries (step-index with varying numerical apertures and core size, GRIN, W-type, H-type and a realistic step-index geometry with a central dip), and different mode-specific intensities ().
+We used 1500 fiber measurements to train the network and 175 for testing.
+3.3.2 Training
 The S-MMF dataset is used to test if the Latent Unmixing method can generalize to other types of distributions; here, additive sinusoidal signals with varying spatial intensities.
@@
 The Latent Unmixing method was trained to map the input data to a latent space of length 60, where 6 bandpass filters of width 10 are applied to separate the space into 6 target modes (Fig.
-7).
+8).
+3.3.3 Results
 Results demonstrate that the Latent Unmixing method can retrieve the spatial distributions of the modes (Fig.
@@
 The total relative intensity is obtained by summing the pixel values over the predicted image of each mode.
-The predicted and ground truth fractions of intensity highly correlate 3 over all numbers of propagating modes.
-Interestingly, our method seems to perform better when the fiber has some intensity 4 or 5 modes than 2 or 3 modes.
+The predicted and ground truth fractions of intensity highly correlate 2 over all numbers of propagating modes.
+Interestingly, our method seems to perform better when the fiber has some intensity in 4 or 5 modes than 2 or 3 modes.
 This goes against the trend of other methods that generally perform poorly for increased numbers of modes [35].
@@
 In this study, we crafted datasets that accurately represent real experiments, while providing an associated ground truth to evaluate the performance of the proposed latent unmixing approach.
-The latent unmixing method also offers interpretability opportunities that were briefly covered in this work.
+The Latent Unmixing method also offers interpretability opportunities.
 By mapping to a pixel latent space with more values than the number of channels in the ground truth (L=28 for four-channel MNIST and FLIM, L=60 for six-channel S-MMF), the network can learn to use the extra values to encode its certainty.
@@
 This is a novel use of deep neural networks, of great interesting for a range of imaging applications, showcasing the capacity of deep networks to process sparse information in its entirety.
-Acknowledgments
-We acknowledge the contributions of Andréanne Deschênes, Antoine Ollier and David Vincent for the scripts used to test the phasor method and maximum likelihood estimation, and Jean-Michel Bellavance for acquiring the STED images.
-This work was supported by the Neuronex Initiative (2014862) - National Science Foundation & Fond de Recherche du Québec - Santé, by the Natural Sciences and Engineering Research Council of Canada - Discovery Grant (DGECR-00200-2019), by the Canada Research Chair Tier 2 (CRC-0126-2019) and Canada-CIFAR AI Chair.
-CB and VB are both supported by NSERC PhD scholarships.
+Acknowledgements We thank Alexy Pelletier-Rioux for the intensity images used for validation, and Albert Michaud-Gagnon and Marie Lafontaine for the implementation of the 2-species phasors unmixing method.
+Declarations
+4.1 Funding
+Funding was provided by grants from the Natural Sciences and Engineering Research Council of Canada (NSERC) (RGPIN-06704-2019 to F.L.C.
+and RGPIN-2019-06706 to C.G.), Fonds de Recherche Nature et Technologie (FRQNT) Team Grant (2021-PR-284335 to F.L.C and C.G.), Sentinel North Initiative funded by Canada First Research Excellence Fund (2020-2024 Major Call for Proposal to F.L.C.
+and C.G.), the Canadian Institute for Health Research (CIHR) (F.L.C.), and the Neuronex Initiative (National Science Foundation 2014862, Fond de recherche du Québec - Santé 295824, ) (F.L.C.).
+C.G.
+is a CIFAR Canada AI Chair and F.L.C.
+is a Canada Research Chair Tier II.
+C.B.
+is supported by scholarships from NSERC, from the Fonds de Recherche Nature et Technologie (FRQNT) Quebec, from the FRQNT strategic cluster UNIQUE, and by a Leadership and Scientific Engagement Award from Université Laval.
+A.D is supported by … J.-M.
+B.
+is supported by ….
+V.B.
+is supported by a scholarship from NSERC.
+4.2 Competing interests
+The authors have no competing interests to declare that are relevant to the content of this article.
+4.3 Code availability
+All the code used to generate the simulated datasets, train the models, and produce the results presented here is available to download at https://anonymous.4open.science/r/LatentUnmixing-101C/.
+All experiments were conducted using Python 3.8.10, with the library versions indicated in the repository.
+4.4
+Consent for publication
+Data availability
+Materials availability
+Author contribution
+If any of the sections are not relevant to your manuscript, please include the heading and write ‘Not applicable’ for that section.
 Supplementary Material
 A. Acquisition of microscopy images
-The STED-FLIM dataset was generated from real 4-colors fluorescence microscopy images acquired with a commercial Abberior STED system.
-Four proteins of interest were marked with fluorescent dyes that have excitation and emission spectral signatures that allow them to be acquired sequentially.
-The neuronal marker MAP2 was tagged with ATTO490LS, the postynaptic marker PSD95 with STAR635, the presynaptic marker VGLUT1 with Alexa488, and the RNA binding protein FUS with Alexa594.
-STAR635 and Alexa594 were imaged with the STED modality (resolution 20 nm), while ATTO490LS and Alexa488 were imaged with the confocal modality (resolution 200 nm).
+Sample preparation
+Neuronal cultures from rats were prepared following established protocols [29] in accordance with procedures approved by the animal care committee of Université Laval.
+Following dissociation from dissected cortexes (hippocampi for FLIM) of P0-P1 postnatal rats, cells were plated onto PDL-Laminin coated glass coverslips (12 mm) in a 24-well plate at a density of 100k cells/well (40k cells/well for FLIM).
+Infection with either a shFUS virus or a PLKO virus was done at 5 of 14 Days in vitro at a concentration of 5 or 10 MOI (4-color only).
+Fixation with 4% PFA was done at 4,7 or 11 days post infection (14 days in vitro for FLIM)
+After fixation, the cells underwent three 5-minute washes in phosphate-buffered saline (PBS) supplemented with 100 mM glycine.
+Subsequently, the cells were permeabilized and blocked with a blocking solution composed of 0.1% Triton X-100 and 2% goat serum (GS) in PBS for 30 minutes before proceeding to immunostaining.
+Immunostaining involved incubation with primary antibodies (PAB) and secondary antibodies (SAB) diluted in the blocking solution at room temperature (see Supp.
+Table 1).
+Primary antibodies were allowed to incubate for 2 hours, followed by three PBS washes.
+Secondary antibodies were then applied for 1 hour and subsequently washed three times with PBS.
+Finally, coverslips were mounted onto slides using Polyvinyl alcohol mounting medium with DABCO (Sigma, cat:10981).
+| Primary antibodies | ||||
+| Antibody | Supplier | Catalogue no. | Dilution | Ref. |
+| Mouse anti-Bassoon | Enzo | ADI-VAM-PS003 | 1 : 500 | [Dani2010, Tang2016, 52] |
+| Rabbit anti-Homer1 | Synaptic Systems | 160003 | 1 : 500 | [Lagache2018, 52] |
+| Mouse anti-MAP2 | Sigma-Aldrich | M4403 | 1 : 1000 | |
+| Rabbit anti-FUS | Sigma-Aldrich | HPA008784 | 1 : 500 | |
+| Guinea Pig anti-VGLUT1 | Sigma-Aldrich | Ab5905 | 1 : 500 | |
+| Fluotag X2 Anti-PSD95 STAR 635P | Nanotag Biotechnologies | Fluotag X2 Anti-PSD95 635P | 1 : 500 | |
+| Secondary Antibodies | ||||
+| Antibody | Company | Catalogue no. | Dilution | |
+| GAM CF594 | Sigma-Aldrich | SAB4600321 | 1 : 500 | |
+| GAR STAR ORANGE | Abberior | STORANGE-1002-500UG | 1 : 250 | |
+| GAM ATTO490LS | Hypermol | 2109-1MG | 1 : 250 | |
+| GAGP Alexa Fluor 488 | Sigma-Aldrich | A-11073 | 1 : 250 | |
+| GAR Alexa Fluor 594 | Invitrogen | A32740 | 1 : 100 | |
+Imaging parameters
+The STED-FLIM dataset was generated from real 4-color fluorescence microscopy images acquired with a commercial Abberior Expert Line STED microscope.
+Four proteins of interest were labelled with fluorescent dyes that have excitation and emission spectral signatures that allow them to be acquired simultaneously using line steps.
+The dendritic marker MAP2 was tagged with ATTO 490LS, the postynaptic marker PSD95 with STAR635P, the presynaptic marker VGLUT1 with Alexa Fluor 488, and the RNA binding protein FUS with Alexa Fluor 594.
+STAR 635P and Alexa Fluor 594 were imaged with the STED modality (resolution 20 nm), while ATTO 490LS and Alexa Fluor 488 were imaged with the confocal modality (resolution 200 nm).
+Further details can be found in Table 2.
+To acquire the real FLIM images, we equipped the microscope with a Time Tagger Ultra 8 from Swabian Instruments GmbH.
+The pre-synaptic protein Bassoon was tagged with CF594 and the post-synaptic protein Homer1 was tagged with STAR Orange.
+These two fluorescent dyes have similar excitation and emission spectra and were imaged using a single imaging channel.
+Further details can be found in Table 3.
+| Parameter | Confocal Alexa Fluor 488 | Confocal Atto 490LS | STED Alexa Fluor 594 | STED STAR 635P |
+| Excitation wavelength | 485 nm | 485 nm | 561 nm | 640 nm |
+| Excitation power | ||||
+| Depletion wavelength | 775 nm | 775 nm | ||
+| Depletion powers | 156 mW | 156 mW | ||
+| Detection wavelengths | 500-550 nm | 650-720 nm | 605-625 nm | 650-720 nm |
+| Pixel dwelltime | ||||
+| Line steps | 2 | 2 | 10 | 6 |
+| Detection gating delay | 750 ps | 750 ps | ||
+| Detection gating width | 8 ns | 8 ns | ||
+| Pixel size | ||||
+| Pinhole size | 1.0 A.U. | 1.0 A.U. | 1.0 A.U. | 1.0 A.U. |
+| Parameter | Confocal FLIM | STED FLIM |
+| Excitation wavelength | 561 nm | 561 nm |
+| Excitation power | ||
+| Depletion wavelength | 775 nm | |
+| Depletion powers | 44, 88, 132, 176 mW | |
+| Detection wavelengths | 605-625 nm | 605-625 nm |
+| Pixel dwelltime | ||
+| Line steps | 4 | 8 |
+| Detection time bins | 250 bins of range of | 250 bins of range of |
+| Pixel size | ||
+| Pinhole size | 1.0 A.U. | 1.0 A.U. |
 B. Architecture and training parameters
@@
 To make sure that the time dimension is exploited by the network to solve the unmixing task, a 2D version of the 3D-MNIST dataset was generated by omitting lines 23-26 and 30-33 in algo.
-1; that is, by simply not creating the time dimension after creating the noisy overlapping digits.
+2; that is, by simply not creating the time dimension after creating the noisy overlapping digits.
 Results obtained by this 2D U-Net are shown in figure 3.
@@
 Maximum likelihood estimation is used to fit the parameters A, B and C in equation 2 for the STED-FLIM test set.
-The confusion matrices between true and predicted photon counts for each channel are shown in Supp. fig. 6(a). The MLE method tends to assign most photons to the shortest lifetime component, A.
+The confusion matrices between true and predicted photon counts for each channel are shown in supp. fig. 6. The MLE method tends to assign most photons to the shortest lifetime component, A.
 We also applied MLE to fit the relative abundance of two components with lifetimes of 2 and 5 ns, which approximately correspond to the lowest and highest lifetime of dyes used for FLIM microscopy.
@@
 fig.
-6(b), with example results from the FLIM-STED test set in Supp.
+7, with example results from the FLIM-STED test set in Supp.
 Fig.
 8 and 9.
-G. Code availability
-All the code used to generate the 3D-MNIST dataset, train the models, and produce the results presented here is available at https://github.com/FLClab/LatentUnmixing.
+G. Unconditional latent unmixing for FLIM
+I. Code availability
+All the code used to generate the 3D-MNIST dataset, train the models, and produce the results presented here is available to download at https://anonymous.4open.science/r/LatentUnmixing-101C/.
 All experiments were conducted using Python 3.8.10, with the library versions indicated in the repository.